How many gallons of a 80% antifreeze solution must be mixed with 80 gallons of 30% antifreeze to get a mixture that is 70% antifreeze
Answer by lwsshak3(11628)

How many gallons of a 80% antifreeze solution must be mixed with 80 gallons of 30% antifreeze to get a mixture that is 70% antifreeze
**
let x=gallons of 80% solution to be mixed
x+80=gallons go final 70% mixture
..
80%*x+30%*80=70%*(x+80)
.8x+24=.7x+56
.1x=32
x=320
ans:
gallons of 80% solution to be mixed=320